# Formation of clusters and coarsening in weakly interacting diffusions

### v1

@@ 2.4. Explicit number of initial clusters for the Hegselmann–Krause prototype
For $w$ as defined in (1.4), we have for all $\ell\in(0,\frac{1}{2})$ and $k\in\mathcal{K}$ that
$\int_{[0,1]}W_{\gamma=1,\ell}(x)\cos(kx)\mathop{}\!\mathup{d}x=\frac{2\cos(k\ell)}{k^{2}}-\frac{2\sin(k\ell)}{k^{3}\ell}.$ (2.15)
One can show that this is minimized for $k=2\pi$ which is the smallest positive element in $\mathcal{K}$. Taylor’s formula around $\ell=0$ yields $\widehat{W}_{1,\ell}(k)=\int_{[0,1]}W_{\gamma=1,\ell}(x)\cos(kx)\mathop{}\!\mathup{d}x=-\frac{2}{3}\ell^{2}+\frac{1}{12}k^{2}\ell^{4}+\dots$, so that
$\gamma_{\sharp}\sim\frac{3}{2\ell^{2}}\qquad\text{as }\ell\to 0.$ (2.16)
In order to determine the number of initial clusters, we note that
@@
$\psi^{\prime}(k)$ $=-2k+\frac{2\gamma}{\ell}\left(\frac{\ell\cos(\ell k)}{k}-\frac{\sin(\ell k)}{k^{2}}\right)+2\gamma\ell\sin(k\ell)$ (2.18)
@@
$\frac{1}{\gamma}=\frac{1}{\ell}\left(\frac{\ell\cos(\ell k)}{k^{2}}-\frac{\sin(\ell k)}{k^{3}}\right)+\ell\frac{\sin(k\ell)}{k}$ (2.19)
@@ 3. Stationary and metastable states
Simulations suggest that for generic initial conditions, the PDE converges to a single-cluster state as $t\to\infty$ unless $\gamma$ is small enough, in which case the PDE converges to the uniform state $\rho_{0}$. This is in line with the fact that the mean-field PDE has a discontinuous phase transition at some $\gamma_{c}$, so that for $\gamma>\gamma_{c}$, there exist global minimizers of the free energy which are non-uniform (see Section 3.1 below for a proof). Moreover, in Theorem 3.3 we prove that all global minimizers are either the uniform state or single-cluster states in the sense that they are symmetrically decreasing. The argument generalizes the strict version of the Riesz rearrangement inequality from [Bec93] to compactly supported kernels, crucially using the analyticity of stationary states which we show in Lemma 3.7. A heuristic argument as well as numerical results as explained in Section 3.3 suggest that for large $\gamma$, the global minimizer of the free energy is in fact approximately (a circular version of) a Gaussian with variance $\sigma^{2}=\frac{\ell}{\gamma w^{\prime\prime}(0)}$.
Figure 3.1. Bifurcation model for the Hegselmann–Krause potential (1.4). Holding $\ell=0.1$ fixed, the bifurcation diagram shows the single-cluster state $\rho_{\gamma}$ as a function of $\gamma$, continuing the single-cluster branch until it ceases to exist. The stationary state $\rho_{\gamma}$ is computed by using Newton’s method to find fixed points of the Kirkwood–Monroe map. The blue curve is the free energy gap $\mathcal{F}_{\gamma,\ell}(\rho_{\gamma})-\mathcal{F}_{\gamma,\ell}(\rho_{\infty})$, where $\rho_{\infty}\equiv 1$ is the uniform stationary state. The red curve shows the $L^{1}$-norm of $\rho_{\gamma}-\rho_{\infty}$. At $\gamma=\gamma_{c}$, there is a discontinuous phase transition. The uniform state $\gamma_{\infty}$ for $\gamma\in[0,\gamma_{\sharp})$ where $\gamma_{\sharp}$ is the point of linear stability. Note that the single-cluster branch exists for $\gamma<\gamma_{c}$ close to $\gamma_{c}$, however, the branch is the global minimizer only for $\gamma>\gamma_{c}$.
@@ 3.1. Existence of a discontinuous phase transition
As shown in [CP10, CGPS20, DGPS23], for small $\gamma$, the uniform state $\rho_{\rm unif}$ is the unique stationary state. Also, for small enough $\ell$, we have a discontinuous phase transition [CGPS20, Def. 5.4] as the following corollary shows. We use the refined result from [GS20, Lemma 5.3], which allows to identify a certain scaling regime between $\gamma$ and $\ell$ in which a discontinuous phase transition is present.
Corollary 3.2.
Assume that $w\colon\mathbb{R}\to\mathbb{R}$ has compact support and that $\int_{\mathbb{R}}w(x)\mathop{}\!\mathup{d}x<0$. Then for sufficiently small $\ell>0$, the potential $W_{\gamma,\ell}$ defined in (1.2) has a discontinuous transition point $\gamma_{c}<\gamma_{\sharp}$. In particular, for $\gamma>\gamma_{c}$ there exists a minimizer $\rho_{\gamma,\ell}$ of ${\mathcal{F}}_{\gamma,\ell}$ distinct from the uniform state $\rho_{\rm unif}(\mathop{}\!\mathup{d}x)=1$
Proof.
The regularity assumption on $w$ ensures that the condition [GS20, (5.2) in Lemma 5.3] is satisfied, where the role of $\epsilon$ is given in terms of $\ell$. The existence of the distinct minimizers of ${\mathcal{F}}_{\gamma,\ell}$ for $\ell$ sufficiently small $\gamma=\gamma_{c}$ is now an immediate consequence of 3.1. ∎
In particular, Corollary 3.2 holds true under 1.1 for $\ell$ sufficiently small and the existence of two energy minimizers for $\gamma=\gamma_{c}$ is illustrated in Fig. 3.1.
3.2. Non-trivial global minimizers of the free energy are symmetrically decreasing
@@
Theorem 3.3 (Global minimizers of the free energy are symmetrically decreasing).
Assume that $w$ satisfies 1.1 and is additionally $C^{2}$ on $(-s_{w},s_{w})$. Fix $\gamma>0$ and $\ell\in(0,1/(2s_{w}))$. Let $\rho$ be a global minimizer of the free energy (1.7). Then there exists a translation $T:\mathbb{T}\to\mathbb{T}$ such that $\rho=\rho^{\ast}\circ T$.
The above Theorem 3.3 provides a non-trivial result if we know that the global minimizer is non-trivial, that is different from the uniform state $\rho_{\rm unif}(\mathop{}\!\mathup{d}x)=1$, which is ensured by the existence of a transition point from Corollary 3.2.
Corollary 3.4 (Single-cluster minimizer).
Assume that $w:\mathbb{R}\to\mathbb{R}$ is as in Theorem 3.3. Then for sufficiently small $\ell>0$ and $\gamma>\gamma_{c}$, the global minimizer $\rho_{\gamma,\ell}$ of ${\mathcal{F}}_{\gamma,\ell}$ is a nontrivial single-cluster state.
@@
From $(\rho^{\sigma})^{\ast}=\rho^{\ast}$ and Riesz’ rearrangement inequality Theorem 3.5, it follows that $I(\rho)\leq I(\rho^{\sigma})\leq I((\rho^{\sigma})^{\ast})=I(\rho^{\ast})$. Since $\rho$ is a minimizer of $I$, it follows that $I(\rho^{\sigma})=I(\rho)$ and the integral on the right-hand side of (3.17) is zero. Since $d(x,\sigma y)<d(x,y)$ for all $x,y\in M^{+}$ and $K$ is a decreasing function, the integrand in (3.17) is non-negative everywhere, therefore it is zero everywhere. Moreover, if $x,y\in M^{+}$ with $d(x,y)<\delta$, we have from $d(x,y)<d(x,\sigma y)$ and (3.5) that
@@
In particular, there exists a constant $c_{1}>0$ such that we have for all smooth $1$-periodic functions $f:\mathbb{R}\to\mathbb{R}$ that
@@
respectively. The free energy (1.7) is then given by the sum $\mathcal{F}(\rho)=S(\rho)+I(\rho)$. The entropy depends only on the masses of the levels sets of $\rho$, so $S(\rho)=S(\rho^{\ast})$. Also note that we are minimizing the free energy and that $W_{\gamma,\ell}(x)$ is increasing as $x$ increases from 0 to $\frac{1}{2}$, so can apply Theorem 3.5 to $I(\rho)$ to obtain $I(\rho)=I(\rho^{\ast})$. By Lemma 3.7, $\rho$ is analytic. From 1.1 we have that $w^{\prime\prime}(0)>0$ which implies (3.5), so by applying Lemma 3.6 to $I(\rho)$, we obtain the claim. ∎
@@ 3.3. Approximate Gaussian steady states
Therefore, a Gaussian with variance $\sigma^{2}=\frac{\ell}{\gamma w^{\prime\prime}(0)}$ satisfies the Kirkwood–Monroe fixed point equation (3.2) approximately for large $\gamma$. Therefore, we expect the single-cluster state to have a variance of $\sigma^{2}$, something which is confirmed by our numerical simulations, see Fig. 3.2. That single-cluster states are approximately Gaussian can also be seen through the particle system by approximating the dynamics of the particles within a cluster by an Orstein–Uhlenbeck process as in [GPY17, Setion 5.3].
@@
Figure 3.2. Single-cluster state are stationary states for $\gamma\gg 1$ (logarithmic $y$-axis). Blue: A numerical steady state of (1.6), computed as fixed point of the Kirkwood–Monroe map (3.2) using Newton’s method. The potential is given as in (1.4) with parameters $\gamma=100$ and $\ell=0.5$. Red: A Gaussian least-squares fit to the numerical steady state, apart from the tails. Its variance is $\sigma^{2}=0.0051$ which approximately equals $\frac{\ell}{\gamma}=0.005$.

### v2

@@ 2.5. Explicit number of initial clusters for the Hegselmann–Krause prototype
For the Hegselmann–Krause prototype (1.4), Lemma 2.1 yields $\gamma_{\sharp}\sim\frac{3}{2}\ell^{-2}$ as $\ell\downarrow 0$. To find the number of initial clusters for (1.4), we calculate $\widehat{W}_{\gamma=1,\ell}(k)=k^{-2}2\cos(k\ell)-k^{-3}\ell^{-1}{2\sin(k\ell)}$, therefore
@@
$\psi^{\prime}(k)$ $=-2k+\frac{2\gamma}{\ell}\left(\frac{\ell\cos(\ell k)}{k}-\frac{\sin(\ell k)}{k^{2}}\right)+2\gamma\ell\sin(k\ell).$ (2.29)
@@
$\frac{1}{\gamma}=\frac{1}{\ell}\left(\frac{\ell\cos(\ell k)}{k^{2}}-\frac{\sin(\ell k)}{k^{3}}\right)+\ell\frac{\sin(k\ell)}{k}.$ (2.30)
@@ 3. Stationary and metastable states
Simulations suggest that for generic initial conditions, the PDE converges to a single-cluster state as $t\to\infty$ unless $\gamma$ is small enough, in which case the PDE converges to the uniform state $\rho_{0}$. This is in line with the fact that the mean-field PDE has a discontinuous phase transition at some $\gamma_{c}$, so that for $\gamma>\gamma_{c}$, there exist global minimisers of the free energy which are non-uniform (see Section 3.1 below for a proof). Moreover, in Theorem 3.4 we prove that all global minimisers are either the uniform state or single-cluster states in the sense that they are symmetrically decreasing. The argument generalizes the strict version of the Riesz rearrangement inequality from [Bec93] to compactly supported kernels, crucially using the analyticity of stationary states which we show in Lemma 3.8. A heuristic argument as well as numerical results as explained in Section 3.3 suggest that for large $\gamma$, the global minimiser of the free energy is in fact approximately (a circular version of) a Gaussian with variance $\sigma^{2}=\frac{\ell}{\gamma w^{\prime\prime}(0)}$.
Figure 3.1. Bifurcation model for the Hegselmann–Krause potential (1.4). Holding $\ell=0.1$ fixed, the bifurcation diagram shows the single-cluster state $\rho_{\gamma}$ as a function of $\gamma$, continuing the single-cluster branch until it ceases to exist. The stationary state $\rho_{\gamma}$ is computed by using Newton’s method to find fixed points of the Kirkwood–Monroe map. The blue curve is the free energy gap $\mathcal{F}_{\gamma,\ell}(\rho_{\gamma})-\mathcal{F}_{\gamma,\ell}(\rho_{\rm unif})$, where $\rho_{\rm unif}\equiv 1$ is the uniform stationary state. The red curve shows the $L^{1}$-norm of $\rho_{\gamma}-\rho_{\rm unif}$. At $\gamma=\gamma_{c}$, there is a discontinuous phase transition. The uniform state $\gamma_{\infty}$ for $\gamma\in[0,\gamma_{\sharp})$ where $\gamma_{\sharp}$ is the point of linear stability. Note that the single-cluster branch exists for $\gamma<\gamma_{c}$ close to $\gamma_{c}$, however, the branch is the global minimiser only for $\gamma>\gamma_{c}$.
@@ 3.1. Existence of a discontinuous phase transition
As shown in [CP10, CGPS20, DGPS23], for small $\gamma$, the uniform state $\rho_{\rm unif}$ is the unique stationary state. Also, for small enough $\ell$, we have a discontinuous phase transition [CGPS20, Def. 5.4] as the following result shows.
Proposition 3.2.
Let $w\colon\mathbb{R}\to\mathbb{R}$ have compact support and moreover assume that $\int_{\mathbb{R}}w(x)\mathop{}\!\mathup{d}x<0$. Then, for all sufficiently small $\ell>0$, the potential $W_{\gamma,\ell}$ defined in (1.2) has a discontinuous transition point $\gamma_{c}<\gamma_{\sharp}$. In particular, for $\gamma>\gamma_{c}$ there exists a global minimiser $\rho_{\gamma,\ell}$ of ${\mathcal{F}}_{\gamma,\ell}$ distinct from the uniform state $\rho_{\rm unif}(\mathop{}\!\mathup{d}x)=1$.
In particular, Proposition 3.2 holds true under 1.1 for $\ell$ sufficiently small. The existence of two energy minimisers for $\gamma=\gamma_{c}$ is illustrated in Fig. 3.1. Below in Section 3.4, we will prove that
$\limsup_{\ell\downarrow 0}\frac{\gamma_{c}}{2\Delta^{-1}\ell^{-1}\log\left(\ell^{-1}\right)}\leq 1,$ (3.4)
where $\Delta=-w(0)>0$. The proof for Proposition 3.2 is based on the following result, which shows that sufficiently localised attractive interacting potentials lead to discontinuous phase transitions [CGPS20, Cor. 5.14].
Proposition 3.3.
[CGPS20, Corollary 5.14] Let $\{W^{(n)}\}_{n\in\mathbb{N}}$ be a sequence of interaction potentials with $\|W^{(n)}\|_{L^{1}}=C>0$ for all $n\in\mathbb{N}$ such that $W^{(n)}\rightarrow-C\delta_{0}$ in the sense of distributions as $n\rightarrow\infty$. Then for $n$ large enough, the associated free energy
$\mathcal{F}_{\gamma}^{(n)}(\rho)=\int_{\mathbb{T}}\rho(x)\log\rho(x)\mathop{}\!\mathup{d}x+\frac{1}{2}\int_{\mathbb{T}\times\mathbb{T}}W^{(n)}(x-y)\rho(x)\rho(y)\mathop{}\!\mathup{d}x\mathop{}\!\mathup{d}y$ (3.5)
possesses a discontinuous transition point at some $\gamma_{c}^{(n)}<\gamma_{\sharp}^{(n)}$, where the point of linear stability $\gamma_{\sharp}^{(n)}$ is defined by replacing $W_{\gamma=1,\ell}$ in the definition  (2.6) by $W^{(n)}$.
Proof of Proposition 3.2.
Let $(\ell_{n})_{n}\in\mathbb{N}$ be a sequence such that $\ell_{n}\downarrow 0$. Then, $W^{(n)}:=W_{\gamma/\ell_{n}^{2},\ \ell_{n}}$ satisfies the requirement of Proposition 3.3. ∎
3.2. Non-trivial global minimisers of the free energy are symmetrically decreasing
@@
Theorem 3.4 (Global minimisers of the free energy are symmetrically decreasing).
Assume that $w$ satisfies 1.1 and is additionally $C^{2}$ on $(-s_{w},s_{w})$. Fix $\gamma>0$ and $\ell\in(0,1/(2s_{w}))$. Let $\rho$ be a global minimiser of the free energy (1.7). Then there exists a translation $T:\mathbb{T}\to\mathbb{T}$ such that $\rho=\rho^{\ast}\circ T$.
The above Theorem 3.4 provides a non-trivial result if we know that the global minimiser is non-trivial, that is different from the uniform state $\rho_{\rm unif}(\mathop{}\!\mathup{d}x)=1$, which is ensured by the existence of a transition point from Proposition 3.2.
Corollary 3.5 (Single-cluster minimiser).
Assume that $w:\mathbb{R}\to\mathbb{R}$ is as in Theorem 3.4. Then for sufficiently small $\ell>0$ and $\gamma>\gamma_{c}$, the global minimiser $\rho_{\gamma,\ell}$ of ${\mathcal{F}}_{\gamma,\ell}$ is a nontrivial single-cluster state.
@@
From $(\rho^{\sigma})^{\ast}=\rho^{\ast}$ and Riesz’ rearrangement inequality Theorem 3.6, it follows that $I(\rho)\leq I(\rho^{\sigma})\leq I((\rho^{\sigma})^{\ast})=I(\rho^{\ast})$. Since $\rho$ is a minimiser of $I$, it follows that $I(\rho^{\sigma})=I(\rho)$ and the integral on the right-hand side of (3.19) is zero. Since $d(x,\sigma y)<d(x,y)$ for all $x,y\in M^{+}$ and $K$ is a decreasing function, the integrand in (3.19) is non-negative everywhere, therefore it is zero everywhere. Moreover, if $x,y\in M^{+}$ with $d(x,y)<\delta$, we have from $d(x,y)<d(x,\sigma y)$ and (3.7) that
@@
In particular, there exists a constant $c_{1}>0$ (depending on $\gamma,\ell$ and $w$) such that we have for all smooth $1$-periodic functions $f:\mathbb{R}\to\mathbb{R}$ that
@@
respectively. The free energy (1.7) is then given by the sum $\mathcal{F}(\rho)=S(\rho)+I(\rho)$. The entropy depends only on the masses of the levels sets of $\rho$, so $S(\rho)=S(\rho^{\ast})$. Also note that we are minimizing the free energy and that $W_{\gamma,\ell}(x)$ is increasing as $x$ increases from 0 to $\frac{1}{2}$, so we can apply Theorem 3.6 to $I(\rho)$ to obtain $I(\rho)\geq I(\rho^{\ast})$. Since $\rho$ is a global minimiser of $\mathcal{F}$, this implies that $I(\rho)=I(\rho^{\ast})$. By Lemma 3.8, $\rho$ is analytic. From 1.1 we have that $w^{\prime\prime}(0)>0$ which implies (3.7), so by applying Lemma 3.7 to $I(\rho)$, we obtain the claim. ∎
@@ 3.3. Approximate Gaussian steady states
Therefore, a Gaussian with variance $\sigma^{2}=\frac{\ell}{\gamma w^{\prime\prime}(0)}$ satisfies the Kirkwood–Monroe fixed point equation (3.2) approximately for large $\gamma$. Therefore, we expect the single-cluster state to have variance $\sigma^{2}$, which is confirmed by our numerical simulations (see Fig. 3.2). That single-cluster states are approximately Gaussian can also be seen through the particle system by approximating the dynamics of the particles within a cluster by an Orstein–Uhlenbeck process as in [GPY17, Setion 5.3].
@@
Figure 3.2. Approximate Gaussian single-cluster states are stationary for $\gamma\gg 1$ (logarithmic $y$-axis). Blue: A numerical steady state of (1.6), computed as fixed point of the Kirkwood–Monroe map (3.2) using Newton’s method. The potential is given as in (1.4) with parameters $\gamma=100$ and $\ell=0.5$. Red: A Gaussian least-squares fit to the numerical steady state. Its variance is $\sigma^{2}=0.0051$ which approximately equals $\frac{\ell}{\gamma}=0.005$.
3.4. Upper bound on the critical interaction strength $\gamma_{c}$.
Next, we give a rigorous upper bound on the critical interaction strength $\gamma_{c}$ by comparing the free energy of the uniform state with the free energy of a rectangular function.
Proposition 3.10.
Assume that $w$ satisfies 1.1 and that $\Lambda>0$ is such that
$w(x)\leq-\Delta+\frac{\Lambda}{2}x^{2}\quad\text{for all }x\in\mathbb{R},$ (3.42)
where $\Delta:=-w(0)>0$. (For example, for (1.4) we can take $\Lambda=1$.) Then, for all $\ell>0$ such that $\ell<\min\left\{\frac{1}{2},\frac{\Delta}{m_{w}},\frac{1}{2s_{w}}\right\}$ where $m_{w}=-\int_{\mathbb{R}}w(x){\rm d}x>0$ , we have that
$\gamma_{c}\leq\gamma_{c+}(\ell):=-\frac{W_{-1}\left(-\lambda(\ell)\ell\exp(c(w))\right)}{\lambda(\ell)},$ (3.43)
where $\lambda(\ell):=\ell\Delta-\ell^{2}m_{w}>0$ and $c(w):=\log\left(\frac{12}{\Lambda}\right)+\frac{1}{2}$ and $W_{-1}$ is the lower branch of the Lambert W function. In particular,
$\limsup_{\ell\downarrow 0}\frac{\gamma_{c}}{2\Delta^{-1}\ell^{-1}\log\left(\ell^{-1}\right)}\leq 1.$ (3.44)
Proof.
For simplicity, we work with the domain $\tfrac{1}{2},\tfrac{1}{2}]$ rather than $[0,1]$. For some fixed $a\in(0,\ell/2)$, we consider the rectangular function $\varrho_{a}:[-\tfrac{1}{2},\tfrac{1}{2}]\to[0,\infty)$ defined by
$\varrho_{a}(x):=\begin{cases}\frac{1}{2a}&\text{if }-a\leq x\leq a,\\ 0&\text{otherwise.}\end{cases}$ (3.45)
We want to compare the free energy of $\varrho_{a}$ with the free energy of the uniform state $\rho_{\rm unif}$. The entropy of $\varrho_{a}$ is given by
$\int_{-\frac{1}{2}}^{\frac{1}{2}}\varrho_{a}(x)\log(\varrho_{a}(x)){\rm d}x=-\log(2a).$ (3.46)
Under the assumption that $a<\frac{\ell}{2}$, we can bound the interaction energy from above using (3.42) as follows:
$\frac{1}{2}\iint_{[-\frac{1}{2},\frac{1}{2}]^{2}}W_{\gamma,\ell}(x-y)\varrho_{a}(x)\varrho_{a}(y)\mathop{}\!\mathup{d}x\mathop{}\!\mathup{d}y=\frac{1}{8a^{2}}\iint_{[-\frac{1}{2},\frac{1}{2}]^{2}}W_{\gamma,\ell}(x-y)\mathop{}\!\mathup{d}x\mathop{}\!\mathup{d}y$ (3.47)
$={}$ $\frac{\gamma\ell}{8a^{2}}\iint_{[-a,a]^{2}}w\left(\frac{x-y}{\ell}\right)\mathop{}\!\mathup{d}x\mathop{}\!\mathup{d}y\leq\frac{\gamma\ell}{8a^{2}}\iint_{[-a,a]^{2}}\left(-\Delta+\frac{\Lambda}{2}\frac{(x-y)^{2}}{\ell^{2}}\right)\mathop{}\!\mathup{d}x\mathop{}\!\mathup{d}y$ (3.48)
$={}$ $-\frac{\gamma\ell\Delta}{2}+\frac{\Lambda}{6}\frac{\gamma}{\ell}a^{2}.$ (3.49)
In particular, we have that
${\mathcal{F}}_{\gamma,\ell}(\varrho_{a})$ $\leq-\log(2a)+\frac{\Lambda}{6}\frac{\gamma}{\ell}a^{2}-\frac{\gamma\ell\Delta}{2}.$ (3.50)
On the other hand, the free energy (1.7) of the uniform state $\rho_{\rm unif}$ is given by
${\mathcal{F}}_{\gamma,\ell}(\rho_{\rm unif})$ $=\frac{1}{2}\int_{[0,1]^{2}}W_{\gamma,\ell}(x-y){\rm d}x{\rm d}y=\frac{1}{2}\int_{[0,1]}W_{\gamma,\ell}(z){\rm d}z$ (3.51)
$=\frac{\gamma\ell}{2}\int_{\mathbb{R}}w\left(\frac{x}{\ell}\right){\rm d}x=\frac{\gamma\ell^{2}}{2}\int_{\mathbb{R}}w(x){\rm d}x=-\frac{\gamma\ell^{2}}{2}m_{w},$ (3.52)
where $m_{w}=-\int_{\mathbb{R}}w(x){\rm d}x>0$ and we used that $W_{\gamma,\ell}$ is 11-periodic and that $\ell s_{w}<\frac{1}{2}$. In particular, we have the upper bound on the free energy gap
${\mathcal{F}}_{\gamma,\ell}(\varrho_{a})-{\mathcal{F}}_{\gamma,\ell}(\rho_{\rm unif})\leq-\log(2a)+\frac{\Lambda}{6}\frac{\gamma}{\ell}a^{2}-\frac{\gamma\ell\Delta}{2}+\frac{\gamma\ell^{2}}{2}m_{w}.$ (3.53)
The minimum of the right-hand side in $a\in(0,\ell/2)$ is given by $a_{0}:=\sqrt{3\ell/(\Lambda\gamma)}$ and we obtain
${\mathcal{F}}_{\gamma,\ell}(\varrho_{a_{0}})-{\mathcal{F}}_{\gamma,\ell}(\rho_{\rm unif})\leq\frac{1}{2}\phi(\gamma,\ell):=\frac{1}{2}\left(-\gamma\lambda(\ell)+\log\gamma-\log\ell-c(w)\right),$ (3.54)
where $\lambda(\ell):=\ell\Delta-\ell^{2}m_{w}>0$ and $c(w):=\log\left(\frac{12}{\Lambda}\right)+\frac{1}{2}$.
Note that $\phi(\gamma,\ell)<0$ is equivalent to
$\exp\left(-\gamma\lambda(\ell)\right)\gamma<\ell\exp(c(w))\Leftrightarrow\exp\left(-\gamma\lambda(\ell)\right)\cdot(-\lambda(\ell)\gamma)<-\lambda(\ell)\ell\exp(c(w)).$ (3.55)
The function $x\mapsto x\exp(x)$ is strictly decreasing on $(-\infty,-1)$ with inverse function $W_{-1}:[-e^{-1},0)\to(-\infty,1]$, the lower real branch of the Lambert $W$-function. Therefore, we have $\phi(\gamma,\ell)<0$ for all $\gamma>\gamma_{c+}(\ell)$, where
$\gamma_{c+}(\ell):=-\frac{W_{-1}\left(-\lambda(\ell)\ell\exp(c(w))\right)}{\lambda(\ell)}>0$ (3.56)
By means of (3.53) and 3.1, this proves (3.43).
We will now prove that
$\lim_{\ell\downarrow 0}\frac{\gamma_{c+}(\ell)}{2\Delta^{-1}\ell^{-1}\log\left(\ell^{-1}\right)}=1,$ (3.57)
which will then imply (3.44). To show (3.57), fix $\ell$ sufficiently small and let $x:=\lambda(\ell)\gamma_{c+}(\ell)$ and $y:=\lambda(\ell)\ell\exp(c(w))$. Then, $x\exp(-x)=y$ and so in particular
$x=\log(y^{-1})+\log(x).$ (3.58)
Assuming $x\geq 1$, (3.58) implies that $x\geq\log(y^{-1})$. On the other hand, we have that
$3\log(y^{-1})\exp(-3\log(y^{-1}))=3\log(y^{-1})y^{3}<3y^{2}<y,$ (3.59)
if $y$ is small enough. Since $t\mapsto t\exp(-t)$ is strictly decreasing on $(-\infty,-1)$, this implies that $x<3\log(y^{-1})$. Plugging this into (3.58), we obtain that $x\leq\log(y^{-1})+\log(\log(y^{-1}))+\log 3$. In particular, we have that
$\frac{W_{-1}(-y)}{\log\left(y^{-1}\right)}=\frac{-x}{\log\left(y^{-1}\right)}\in\left(1,1+\frac{\log(\log(y^{-1}))+\log 3}{\log(y^{-1})}\right)$ (3.60)
which proves (3.57). ∎
Remark 3.11.
Instead of approximating the single-cluster stationary states of the PDE with rectangular functions, one could alternatively use analogues of the Gaussian on the circle, such as the von Mises distribution or the wrapped Gaussian. While these choices would provide a more accurate representation of the cluster shape, the resulting upper bound for $\gamma_{c+}$ would remain of the same order in $\ell$, and the calculations would become more technical.